An apartment building contains 12 units consisting of one- and two-bedroom apartments that rent for $360 and $450 per month, respectively. When all units are rented, the total monthly rental is $4,950. What is the number of two-bedroom apartments?
CAN I GET SOME HELP?

Answer by mananth(16948)

let two bed room be x ... Rent $450
one bed room be 12-x..... rent $360
Total there are 12
..
Total rent recieved = $4950
450x+360(12-x)=4950
450x +4320-360x =4950
90x+4320=4950
90x=630
x= 630/90
x=7 two bed room apartments